# TACS-Graphs: Traversability-Aware Consistent Scene Graphs for Ground Robot Indoor Localization and Mapping

### v1

@@ II.C Traversability Mapping
Traversability mapping is essential for the safe exploration and efficient navigation of autonomous ground robots. It supports key robotic functions such as obstacle recognition [20, 21, 22], static map generation [23, 24], odometry estimation [25, 26], and global localization [27].
@@ IV.C2 Quantitative Evaluation
$N_{\mathrm{room}}$ $\sigma_{\mathrm{room}}$downarrow DCS(%) $\uparrow$ $t_{\mathrm{FGO}}$(s) $\downarrow$
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 1.0 0.0 0.0 0.02
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 8.0 3.3 6.2 0.18
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 10.0 2.1 3.2 0.93
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 14.3 2.1 14.2 0.28
@@ IV.D1 Loop Closure Performance
ATE (m) $\downarrow$ $N_{\mathrm{loop}}$ ${PR}_{\mathrm{loop}}$uparrow $t_{\mathrm{LD}}$ (s) $\downarrow$
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 0.12 21 0.28 2.26
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 23.06 56.6 0.29 14.45
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 0.42 137 0.35 23.71
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$=$5\mathrm{m}$ 2.55 108.0 0.08 21.15
@@ V Conclusion
Existing 3D scene graph generation methods struggle with room segmentation inconsistencies, limiting their applicability in complex indoor environments. To address this, we proposed TACS-Graphs, which incorporates a traversability map for ground robots to enhance scene graph consistency by mitigating under- and over-segmentation. Additionally, a novel loop closure detection module improves PGO efficiency and accuracy by leveraging scene graph consistency. Our key contribution is the identification of scene graph consistency as a critical challenge and its mitigation through traversability-based room segmentation. Extensive experiments in diverse environments validated our approach. However, as the current approach uses the Manhattan world assumption, future research will extend scene graphs to relax this assumption by incorporating structure-invariant models such as GNNs with traversability maps. This will facilitate adaptive room segmentation, further improving robotic task performance in diverse and challenging environments.

### v2

# TACS-Graphs: Traversability-Aware Consistent Scene Graphs for Ground Robot Localization and Mapping

@@ II.C Traversability Mapping
Traversability mapping is essential for the safe exploration and efficient navigation of autonomous ground robots. It supports key robotic functions such as obstacle recognition [20, 21], static map generation [22, 23], odometry estimation [24, 25], and global localization [26].
@@ IV.C2 Quantitative Evaluation
$N_{\mathrm{room}}$ $\sigma_{\mathrm{room}}$ $\downarrow$ DCS(%) $\uparrow$ $t_{\mathrm{FGO}}$(s) $\downarrow$
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 1.0 0.0 0.0 0.02
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 8.0 3.3 6.2 0.18
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 10.0 2.1 3.2 0.93
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 14.3 2.1 14.2 0.28
@@ IV.D1 Loop Closure Performance
ATE (m) $\downarrow$ $N_{\mathrm{loop}}$ ${PR}_{\mathrm{loop}}$ $\uparrow$ $t_{\mathrm{LD}}$ (s) $\downarrow$
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 0.12 21 0.28 2.26
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 23.06 56.6 0.29 14.45
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 0.42 137 0.35 23.71
@@
S-Graphs+ with $d^{\text{loop}}_{\text{th}}$$5\mathrm{m}$ 2.55 108.0 0.08 21.15
@@ V Conclusion
Existing 3D scene graph generation methods struggle with room segmentation inconsistencies, limiting their applicability in complex indoor environments. To address this, we proposed TACS-Graphs, which incorporates a traversability map for ground robots to enhance scene graph consistency by mitigating under- and over-segmentation. Additionally, a novel loop closure detection module improves PGO efficiency and accuracy by leveraging scene graph consistency. Our key contribution is the identification of scene graph consistency as a critical challenge and its mitigation through traversability-based room segmentation. Extensive experiments in diverse environments validated our approach. However, the current approach is limited by the Manhattan world assumption, which restricts its applicability to rooms with rectilinear structures. As a result, it cannot be effectively applied to environments with irregular or non-orthogonal layouts. Future research will aim to develop more generalizable mapping methods that can handle a wide variety of indoor and outdoor environments, without relying on such structural assumptions.